If A=10ax-4ay+6az and B=2ax+ay, find a the component of A along ay b the magnitude of 3A-B, c a unit vector along A+2B
1
Expert's answer
2020-10-19T13:24:43-0400

"A=10a_x-4a_y+6a_z" and "B=2a_x+a_y"

a) If "A=A_xa_x+A_ya_y+A_za_z" , then the component of "A" along "a_y" is "A_y" .

For "A=10a_x-4a_y+6a_z" we have that "A_y=-4" is the component of "A" along "a_y".


b) "3A-B=3 (10a_x-4a_y+6a_z )-(2a_x+a_y)=(3\times 10-2)a_x+(3\times(-4)-1)a_y+3\times6a_z"

"3A-B=28a_x-13a_y+18a_z"

"|3A-B|=\sqrt{28^2+(-13)^2+18^2}=\sqrt{1277}=35.74"


c) "A+2B= 10a_x-4a_y+6a_z+2(2a_x+a_y)=(10+2\times2)a_x+(-4+2)a_y+6a_z"

"A+2B=14a_x-2a_y+6a_z"

A unit vector along "A+2B" is vector "\frac{14a_x-2a_y+6a_z}{\sqrt{14^2+(-2)^2+6^2}}= \frac{14a_x-2a_y+6a_z}{\sqrt{236}}\approx0.91a_x-0.13a_y+0.39a_z"


Answer: a) "-4" ; b) "35.74" ; c) "0.91a_x-0.13a_y+0.39a_z"